Not sure how to find the slope of the line joining (3, 12) and (15, -2)as a fraction in the simplest form.
Answer by Maths68(1474)

(3, 12) and (15, -2)
Slope of line = Change in y / change in x
m=y2-y1/x2-x1
m=-2-12/15-3
m=-14/12
m=-7/6
Slope = -7/6